A 3550 mL sample of gas at 20.0 degrees C and a pressure of 650.0 torr is allowed to expand to a volume of 8.00 L. The final pressure of the gas is 32.0 kPa. What is the final temp. in degrees celsius?
1
Expert's answer
2013-05-17T10:15:47-0400
As the pressure goes up, the temperature also goes up, and vice-versa. Initial and final volumes and temperatures under constant pressure can be calculated.

P1V1/T1 = P2V2/T2 = P3V3/T3 etc.

86.66 kPa *3,55 L/ 293 = 32.0 kPa*8.00 L/T2

1,049 = 32.0 kPa*8.00 L/T2
T2 = 32.0 kPa*8.00 L / 1.049
T2 = 244.041 K = - 28.95 degrees C
Answer: T2 = 244.041 K